OpenStudy (anonymous):

Use graphs and tables to find the limit and identify any vertical asymptotes of the function. limit of 1 divided by the quantity x minus 8 as x approaches 8 from the left

myininaya (myininaya):

\[\lim_{x \rightarrow 8^-} \frac{1}{x-8}\]

myininaya (myininaya):

first of all does y=1/(x-8) have any vertical asymptotoes?

OpenStudy (anonymous):

no it doesn't

myininaya (myininaya):

When you have zero on bottom, you either have a hole or a vertical asymptote. This graph has no holes because their is no common factor (x-8) on top and bottom. So you have a vertical asmptote when x-8=0.

OpenStudy (anonymous):

oh okay, i get that part now

OpenStudy (anonymous):

so how would i find the asmptote

myininaya (myininaya):

By finding when the bottom is 0.

myininaya (myininaya):

Did you read my paragraph up there? "So you have a vertical asymptote when x-8=0."

myininaya (myininaya):

You need to solve x-8=0 for x

OpenStudy (anonymous):

x = 8